In a class of 10 students 4 are girls 3 are selected at random.find the probability
(a) 2 girls are selected
(b) atleast 2 boys are selected
(c) atmost 2 girls are selected

Answer by Boreal(15235)

10C3=120 ways to choose
2 girls and 1 boy: numerator is 6C1*4C2=6*6=36, so probability is 0.3
at least 2 boys are selected. For 2: 6C2*4C1=60 ways or 0.5 prob.
3 boys would have 6C3*4C0 ways or 20 , so probability is 0.67 for at least 2
at most 2 girls are selected are 2 boys 1 girl (0.5 prob. as stated above)
the last (not asked for is) 3 girls and that is 4 ways or (0.0333 prob.
----
prob. 0 boys is 4/120
1 boy is 36/120
2 boys is 60/120
3 boys is 20/120
They add to 1.